AR3CMONITOR(1) UserCommands AR3CMONITOR(1) NAME ar3cmonitor−AR-3cvisualizationsystemclient SYNOPSIS ar3cmonitor[options][host][port] DESCRIPTION Thear3cmonitorisanAR-3cvisualizationsystemclient. OPTIONS −−acacheopt enableAnalogChacheoptimizations(EXPERIMENTAL) −−conf−locale setconfigurationlanguage −−encoding setconfigencoding −−fullscreen fullscreenmode −h,−−help showthishelp −−lang setuserinterfacelanguage −−maximize showmaximized −−memlowhigh setlowandhighmemorymark −m,−−menu showmenubar −n,−−scheme usescheme(0bydefault,−1forwithoutschemeinit) −−off−update−ival changeofflinemodeupdateinterval(ms) −−rlimitlimit setrequestlimit(experimental,default1) −−update−ival changeupdateinterval(ms) −V,−−version showversion EXAMPLES Connecttoar3cctrlat192.168.2.100ondefaultport: ar3cmonitor192.168.2.100 UseplinterfacelanguageandRussianconfigurationlocale: ar3cmonitor--langpl--conf-localeruRU192.168.2.100 ar3cmonitor March2008 1 AR3CMONITOR(1) UserCommands AR3CMONITOR(1) AUTHOR KrzysztofMazur<krzysiek@podlesie.net> SEEALSO ar3cctrl(1) ar3cmonitor March2008 2